# Theory Uncertainties in the Extraction of $\alpha_{s}$ from Drell-Yan at Small Transverse Momentum

### v1

@@ 3.4.1 General discussion
where $y_{i}$ represent the data, $\lambda_{i}$ the corresponding theory model, and $C_{ij}$ the covariance matrix encoding all experimental uncertainties and correlations between the data points. The $\lambda_{i}$ implicitly depend upon $\alpha_{s}(m_{Z})$ and the parameters of the underlying theory prediction, i.e. the TNPs and nonperturbative parameters discussed in sections 3.2.3 and 3.3. The last term imposes the Gaussian theory constraint on the TNPs, $\theta_{i}$, about their central values of 00 with a prior uncertainty of $\Delta\theta_{i}$, as discussed in section 3.2.3.
@@ 3.4.2 Fit setup
Regarding the PDF dependence, we always use the MSHT20aN3LO PDF set [54], and account for the interplay between $\alpha_{s}(m_{Z})$ and the PDFs in the $q_{T}$ spectrum. Since $\alpha_{s}$ and the PDFs are strongly correlated, we ensure that the PDF set entering the prediction matches the value of $\alpha_{s}(m_{Z})$ being tested: For $\alpha_{s}(m_{Z})<0.118$ we use the set with $\alpha_{s}(m_{Z})=0.117$, and for $\alpha_{s}(m_{Z})>0.118$ we use the set with $\alpha_{s}(m_{Z})=0.119$, appropriately interpolating for intermediate values.
@@ 4.1 Scale variations
Figure -958: Left panel: Relative differences in the $q_{T}$ spectrum of individual scale variations at N4LL. Different colored lines show different classes of variations as defined in section 3.2.2. The red dashed lines show the relative difference from varying $\alpha_{s}(m_{Z})$ for comparison. Right panel: The resulting $\alpha_{s}(m_{Z})$ values when performing the Asimov fit to the central N4LL prediction for each scale variation. Different colors correspond to the same classes of variations as on the left. The grey band shows the fit uncertainty for comparison.
@@
In the left panel of figure -958, we show the breakdown of all scale variations for the $q_{T}$ spectrum at N4LL from SCETlib as discussed in section 3.2.1. Overlaid on these variations are the central predictions for $\alpha_{s}(m_{Z})=0.117$ and $\alpha_{s}(m_{Z})=0.119$, illustrating that the sensitivity to $\alpha_{s}(m_{Z})$ in the $q_{T}$ spectrum is primarily a shape effect. To illustrate the scale-variation scanning in practice, we use the Asimov fit setup described in section 3.4. Our pseudodata corresponds to the N4LL central scale prediction at $\alpha_{s}(m_{Z})=0.118$. We then repeat the fit of $\alpha_{s}(m_{Z})$ for each scale variation, each representing a different theory model. The right panel of figure -958 shows the results, where each point represents the fitted value of $\alpha_{s}(m_{Z})$ for one of the scale variations on the left. Points marked with a cross indicate cases where $\chi^{2}/{\rm n_{\rm dof}}\geq 1.5$, signifying a very poor fit to the pseudodata. The fitted $\alpha_{s}(m_{Z})$ values strongly depend on the choice of scale variation. While some variations have almost no impact on $\alpha_{s}(m_{Z})$, yielding values very close to the nominal value of $0.118$, others yield very different values. This demonstrates that as anticipated the impact on $\alpha_{s}(m_{Z})$ strongly depends on the precise shape of any given variation. As a result, the precise choice of scale variations directly dictates the size of the would-be perturbative uncertainty, making this approach unsuitable for a reliable uncertainty estimate.
@@
$\Delta_{\rm FO}$ $=1.55\times 10^{-3}\,,\qquad$ $\Delta_{f}$ $=0.71\times 10^{-3}\,,$
$\Delta_{\rm resum}$ $=1.73\times 10^{-3}\,,\qquad$ $\Delta_{\rm match}$ $=0.06\times 10^{-3}\,.$ (4.1)
@@
[…]
@@ 4.2.1 Scanning
Figure -957: Left panel: Relative uncertainties in the $q_{T}$ spectrum with TNPs at N3+1LL. The different lines show the impact of varying the corresponding TNP by $+1$ or $-1$, corresponding to 68% theory CL. The yellow band shows their sum in quadrature. Right panel: Corresponding uncertainty on $\alpha_{s}(m_{Z})$ from performing the Asimov fit to the central N3+1LL prediction for each TNP variation. The grey band shows the fit uncertainty for comparison.
@@
The left panel of figure -957 presents the breakdown of the individual TNP variations for the $q_{T}$ spectrum at N3+1LL accuracy. Each TNP is varied by $\Delta\theta_{n}=1$. Since the resulting variations in the spectrum are nearly symmetric, for clarity, we display only the up variations for the anomalous dimensions, hard, and soft functions, and the down variations for the beam functions. Each TNP corresponds to an independent source of uncertainty and is thus varied independently, capturing the correct correlations across the $q_{T}$ spectrum. We can now sum in quadrature the seven independent TNP variations to obtain the total uncertainty band, depicted in yellow.
To perform the scanning, we use the central N3+1LL prediction as Asimov data and repeat the fit for N3+1LL theory models with each TNP varied separately up and down by $\Delta\theta_{n}=1$. The right panel of figure -957 shows the results, where the error bars represent the difference between the fitted $\alpha_{s}(m_{Z})$ and the nominal expected value of $0.118$, i.e., the uncertainty on the fitted $\alpha_{s}(m_{Z})$ due to a given TNP. As for the spectrum, we can now sum the individual uncertainties in quadrature to obtain the total uncertainty on $\alpha_{s}(m_{Z})$, which yields
@@ 4.2.2 Profiling
Figure -956: Left panel: Relative uncertainties in the $q_{T}$ spectrum at N3+1LL before and after profiling the TNPs. The different lines show the post-fit relative impact of each TNP and the orange band the total post-fit uncertainty. The yellow band shows the pre-fit uncertainty corresponding to the yellow band in figure -957. Right panel: Post-fit constraints on the TNPs (error bars) and their impact on $\alpha_{s}(m_{Z})$, with the solid (dashed) grey band showing the impact of the post-fit downward (upward) TNP variations.
@@
The left panel of figure -956 presents the post-fit prediction for the $q_{T}$ spectrum with a breakdown of the post-fit uncertainties, showing the upward variation for each TNP, except for the beam functions which display the downward variation. As a general comment, which applies to this and all analogous plots in the following, note that the individual TNP post-fit variations do not necessarily have to lie within the total post-fit uncertainty, as there are now nontrivial correlations between the TNPs imposed by the data. In particular, if two TNPs are negatively correlated, their combined uncertainty on the spectrum and $\alpha_{s}(m_{Z})$ can be smaller than their individual impacts.
After profiling the TNPs, the uncertainties for both $\alpha_{s}(m_{Z})$ and the $q_{T}$ spectrum are significantly smaller compared to simply scanning the TNPs and the corresponding pre-fit uncertainties. However, focusing solely on the final uncertainty after profiling can also be misleading and may lead to incorrect conclusions. An important aspect when profiling systematic uncertainties is to examine to what extent the data actually pulls and constraints the nuisance parameters in order to ensure that the constrained post-fit theory model is still adequate and allowed. In our case, we have to ensure that the theory model is not overconstrained to the point that neglected yet higher-order terms become relevant for determining the theory uncertainty [34]. The right panel of figure -956 presents a pull/impact plot for all TNPs. The error bars refer to the bottom $x$-axis and show the post-fit constraints for each TNP. The solid (dashed) grey bars refer to the top $x$-axis and show the corresponding impact on $\alpha_{s}(m_{Z})$ due to each TNPs downward (upward) post-fit uncertainty. [^5] The impacts are evaluated by repeating the fit while fixing each TNP one at a time to its post-fit value plus/minus its post-fit uncertainty. The difference of the obtained $\alpha_{s}(m_{Z})$ to the result of the nominal fit gives the impact on $\alpha_{s}(m_{Z})$. The post-fit central value for all TNPs is 00, in perfect agreement with their values in the pseudodata, as should be the case here. The post-fit constraints on the individual TNPs are overall not strong. The individual impacts provide further insight into which TNPs are most important (after profiling) for determining $\alpha_{s}(m_{Z})$. Here, $\gamma_{\nu}$, $S$, and $B_{qg}$ have the biggest post-fit impact on $\alpha_{s}(m_{Z})$ and they are also the ones that are somewhat constrained. This indicates that these parameters have the strongest correlation with $\alpha_{s}(m_{Z})$. Of course, to study the exact correlations between all parameters in detail, one has to examine the post-fit covariance matrix.
@@ 4.2.3 Profiling against N4LL
Figure -955: Left panel: Uncertainties in the $q_{T}$ spectrum at N2+1LL relative to the N4LL, before (yellow band) and after (orange band) profiling the TNPs. The different lines show the post-fit relative impact of each TNP. Right panel: Post-fit constraints on the TNPs (error bars) and their impact on $\alpha_{s}(m_{Z})$, with the solid (dashed) grey band showing the impact of the post-fit downward (upward) TNP variations. The stars indicate the true values of the TNPs.
@@
In figure -955, we consider the case where the theory model is given by the N2+1LL prediction. The left panel shows the pre-fit and post-fit uncertainties for the $q_{T}$ spectrum analogous to the left panel of figure -956 but now relative to the true N4LL result. The profiling again leads to a significant reduction in the uncertainty. Furthermore, we see that the post-fit prediction is now pulled toward the true result, with a significant adjustment of the shape of the spectrum. Whilst the shape of the pre-fit prediction differs from the true shape, the post-fit prediction closely reproduces it.
@@
The right panel of figure -955 shows the corresponding pull plot for the TNPs (analogous to the right panel of figure -956) with the true values of the TNPs marked by stars. The fitted TNPs at N2+1LL now exhibit strong pulls toward their true values, showing that, as for the $q_{T}$ spectrum, the fit is moving the theory model in the right direction also at the level of individual TNPs. This is exactly what we would like to happen in this situation where the data has a higher precision than the pre-fit theory constraint. However, some of the TNPs, such as $\gamma_{\nu}$, $B_{qq}$, and in particular $B_{qg}$, are quite strongly constrained now. A tight post-fit constraint implies that the corresponding parameter at the next higher-order becomes relevant. In other words, we are starting to “overfit” the N2+1LL theory model. The precision of the data is starting to exceed the inherent precision of the theory model, suggesting that it is becoming insufficient and using a higher-order theory-model might be warranted.
@@
Figure -954: Left panel: Uncertainties in the $q_{T}$ spectrum at N3+1LL relative to the N4LL, before (yellow band) and after (orange band) profiling the TNPs. The different lines show the post-fit relative impact of each TNP. Right panel: Post-fit constraints on the TNPs (error bars) and their impact on $\alpha_{s}(m_{Z})$, with the solid (dashed) grey band showing the impact of the post-fit downward (upward) TNP variations. The stars indicate the true values of the TNPs.
@@
In figure -954 we repeat the same exercise with the N3+1LL theory model. The post-fit uncertainties for the $q_{T}$ spectrum are very similar to those in figure -956 when profiling N3+1LL against itself, except that now the central value is shifted toward the true result. The same holds at the level of individual TNPs. Their post-fit uncertainties are very similar to before, while their central values are shifted toward their true values, which is again as we would like it to be. Compared to the N2+1LL case, the TNPs are now less constrained, as expected, with no indications yet of overfitting the theory model.
@@
Figure -953: Perturbative uncertainties on $\alpha_{s}(m_{Z})$ when profiling the TNPs at different orders against the central N4LL prediction. The numerically displayed uncertainties are in units of $10^{-3}$.
@@
In figure -953, we present the results for $\alpha_{s}(m_{Z})$ when profiling the TNPs at various orders against the N4LL pseudodata. The uncertainty on $\alpha_{s}(m_{Z})$ is only slightly larger at N2+1LL than at N3+1LL, which highlights the fact that after profiling it is largely driven by the constraints imposed by the measurement on the theory model. In addition, we also include results at N3+0LL and N4+0LL – these orders serve as approximations of N3+1LL and N4+1LL, respectively, see appendix B.2 for their definition and a more detailed discussion. While these approximations lack the correct formal structure of their corresponding higher orders, we include them here for completeness. Note that although N4+0LL is technically our currently highest available order, we prefer to use N3+1LL as our nominal theory model, because it has the correct theory correlation structure. Furthermore, its theory precision is sufficiently high compared to the measurement precision and after profiling it yields uncertainties comparable to N4+0LL.
@@
At both N2+1LL and N3+1LL we observe a small bias in the fitted central value of $\alpha_{s}(m_{Z})$. First note that when fitting the lower-order theory model against the higher-order pseudodata, we intentionally introduce a mismatch. Without profiling, this naturally leads to a potential bias in $\alpha_{s}(m_{Z})$ as it potentially absorbs some of the deficiencies of the theory model. This type of bias is actually what is estimated and supposed to be covered by the scanning approach (for either TNPs or scale variations). Indeed, fitting the central N2+1LL and N3+1LL models against N4LL, we find central values of $\alpha_{s}(m_{Z})=0.1211$ and $\alpha_{s}(m_{Z})=0.1195$, respectively. These biases compared to the correct value of $0.118$ are indeed covered by the corresponding scanning uncertainties.

### v2

@@ 3.4.1 General discussion
where $y_{i}$ represent the data, $\lambda_{i}$ the corresponding theory model, and $C_{ij}$ the covariance matrix encoding all experimental uncertainties and correlations between the data points. The $\lambda_{i}$ implicitly depend upon $\alpha_{s}(m_{Z})$ and the parameters of the underlying theory prediction, i.e. the TNPs and nonperturbative parameters discussed in sections 3.2.3 and 3.3. Within the fit, $\alpha_{s}(m_{Z})$, the TNPs, and nonperturbative parameters are treated as continuous parameters. The last term imposes the Gaussian theory constraint on the TNPs, $\theta_{i}$, about their central values of 0 with a prior uncertainty of $\Delta\theta_{i}$, as discussed in section 3.2.3.
@@ 3.4.2 Fit setup
Regarding the PDF dependence, we always use the MSHT20aN3LO PDF set [59], and account for the interplay between $\alpha_{s}(m_{Z})$ and the PDFs in the $q_{T}$ spectrum. Since $\alpha_{s}$ and the PDFs are strongly correlated, we ensure that the PDF set entering the prediction is consistent with the value of $\alpha_{s}(m_{Z})$ being tested. For this purpose, we perform a linear interpolation between the available PDF sets for $\alpha_{s}(m_{Z})=0.118$ and $\alpha_{s}(m_{Z})=0.117$ when $\alpha_{s}(m_{Z})<0.118$ and between the PDF sets for $\alpha_{s}(m_{Z})=0.118$ and $\alpha_{s}(m_{Z})=0.119$ when $\alpha_{s}(m_{Z})>0.118$. This interpolation is only used for the PDFs, while the dependence on $\alpha_{s}(m_{Z})$ in the perturbative prediction itself is always treated exactly.
@@ 4.1 Scale variations
Figure 2: Left panel: Relative differences in the $q_{T}$ spectrum of individual scale variations at N4LL. Different colored lines show different classes of variations as defined in section 3.2.2. The red dashed lines show the relative difference from varying $\alpha_{s}(m_{Z})$ for comparison. Right panel: The resulting $\alpha_{s}(m_{Z})$ values when performing the Asimov fit to the central N4LL prediction for each scale variation. Different colors correspond to the same classes of variations as on the left. The grey band shows the fit uncertainty for comparison.
@@
In the left panel of figure 2, we show the breakdown of all scale variations for the $q_{T}$ spectrum at N4LL from SCETlib as discussed in section 3.2.1. Overlaid on these variations are the central predictions for $\alpha_{s}(m_{Z})=0.117$ and $\alpha_{s}(m_{Z})=0.119$, illustrating that the sensitivity to $\alpha_{s}(m_{Z})$ in the $q_{T}$ spectrum is primarily a shape effect. To illustrate the scale-variation scanning in practice, we use the Asimov fit setup described in section 3.4. Our pseudodata corresponds to the N4LL central scale prediction at $\alpha_{s}(m_{Z})=0.118$. We then repeat the fit of $\alpha_{s}(m_{Z})$ for each scale variation, each representing a different theory model. The right panel of figure 2 shows the results, where each point represents the fitted value of $\alpha_{s}(m_{Z})$ for one of the scale variations on the left. Points marked with a cross indicate cases where $\chi^{2}/{\rm n_{\rm dof}}\geq 1.5$, signifying a very poor fit to the pseudodata. The fitted $\alpha_{s}(m_{Z})$ values strongly depend on the choice of scale variation. While some variations have almost no impact on $\alpha_{s}(m_{Z})$, yielding values very close to the nominal value of $0.118$, others yield very different values. This demonstrates that as anticipated the impact on $\alpha_{s}(m_{Z})$ strongly depends on the precise shape of any given variation. As a result, the precise choice of scale variations directly dictates the size of the would-be perturbative uncertainty, making this approach unsuitable for a reliable uncertainty estimate.
@@
$\Delta_{\rm FO}$ $=1.55\times 10^{-3}\,$$\qquad$ $\Delta_{f}$ $=0.71\times 10^{-3}\,$,
$\Delta_{\rm resum}$ $=1.73\times 10^{-3}\,$$\qquad$ $\Delta_{\rm match}$ $=0.06\times 10^{-3}\,.$ (4.1)
@@
$\text{sum of envelopes:}\qquad\Delta_{\rm pert}$ $=\sqrt{\Delta_{\rm FO}^{2}+\Delta_{f}^{2}+\Delta_{\rm resum}^{2}+\Delta_{\rm match}^{2}}=2.43\times 10^{-3}\,$,
$\text{total envelope:}\qquad\Delta_{\rm pert}$ $=1.73\times 10^{-3}\,.$ (4.2)
@@ 4.2.1 Scanning
Figure 3: Left panel: Relative uncertainties in the $q_{T}$ spectrum with TNPs at N3+1LL. The different lines show the impact of varying the corresponding TNP by +1+1 or 1-1, corresponding to 68% theory CL. The yellow band shows their sum in quadrature. Right panel: Corresponding uncertainty on $\alpha_{s}(m_{Z})$ from performing the Asimov fit to the central N3+1LL prediction for each TNP variation. The grey band shows the fit uncertainty for comparison.
@@
The left panel of figure 3 presents the breakdown of the individual TNP variations for the $q_{T}$ spectrum at N3+1LL accuracy. Each TNP is varied by $\Delta\theta_{n}=1$. Since the resulting variations in the spectrum are nearly symmetric, for clarity, we display only the up variations for the anomalous dimensions, hard, and soft functions, and the down variations for the beam functions. Each TNP corresponds to an independent source of uncertainty and is thus varied independently, capturing the correct correlations across the $q_{T}$ spectrum. We can now sum in quadrature the seven independent TNP variations to obtain the total uncertainty band, depicted in yellow.
To perform the scanning, we use the central N3+1LL prediction as Asimov data and repeat the fit for N3+1LL theory models with each TNP varied separately up and down by $\Delta\theta_{n}=1$. The right panel of figure 3 shows the results, where the error bars represent the difference between the fitted $\alpha_{s}(m_{Z})$ and the nominal expected value of $0.118$, i.e., the uncertainty on the fitted $\alpha_{s}(m_{Z})$ due to a given TNP. As for the spectrum, we can now sum the individual uncertainties in quadrature to obtain the total uncertainty on $\alpha_{s}(m_{Z})$, which yields
@@ 4.2.2 Profiling
Figure 4: Left panel: Relative uncertainties in the $q_{T}$ spectrum at N3+1LL before and after profiling the TNPs. The different lines show the post-fit relative impact of each TNP and the orange band the total post-fit uncertainty. The yellow band shows the pre-fit uncertainty corresponding to the yellow band in figure 3. Right panel: Post-fit constraints on the TNPs (error bars) and their impact on $\alpha_{s}(m_{Z})$, with the solid (dashed) grey band showing the impact of the post-fit downward (upward) TNP variations.
@@
The left panel of figure 4 presents the post-fit prediction for the $q_{T}$ spectrum with a breakdown of the post-fit uncertainties, showing the upward variation for each TNP, except for the beam functions which display the downward variation. As a general comment, which applies to this and all analogous plots in the following, note that the individual TNP post-fit variations do not necessarily have to lie within the total post-fit uncertainty, as there are now nontrivial correlations between the TNPs imposed by the data. In particular, if two TNPs are negatively correlated, their combined uncertainty on the spectrum and $\alpha_{s}(m_{Z})$ can be smaller than their individual impacts.
After profiling the TNPs, the uncertainties for both $\alpha_{s}(m_{Z})$ and the $q_{T}$ spectrum are significantly smaller compared to simply scanning the TNPs and the corresponding pre-fit uncertainties. However, focusing solely on the final uncertainty after profiling can also be misleading and may lead to incorrect conclusions. An important aspect when profiling systematic uncertainties is to examine to what extent the data actually pulls and constraints the nuisance parameters in order to ensure that the constrained post-fit theory model is still adequate and allowed. In our case, we have to ensure that the theory model is not overconstrained to the point that neglected yet higher-order terms become relevant for determining the theory uncertainty [34]. The right panel of figure 4 presents a pull/impact plot for all TNPs. The error bars refer to the bottom $x$-axis and show the post-fit constraints for each TNP. The solid (dashed) grey bars refer to the top $x$-axis and show the corresponding impact on $\alpha_{s}(m_{Z})$ due to each TNPs downward (upward) post-fit uncertainty. [^5] The impacts are evaluated by repeating the fit while fixing each TNP one at a time to its post-fit value plus/minus its post-fit uncertainty. The difference of the obtained $\alpha_{s}(m_{Z})$ to the result of the nominal fit gives the impact on $\alpha_{s}(m_{Z})$. The post-fit central value for all TNPs is 0, in perfect agreement with their values in the pseudodata, as should be the case here. The post-fit constraints on the individual TNPs are overall not strong. The individual impacts provide further insight into which TNPs are most important (after profiling) for determining $\alpha_{s}(m_{Z})$. Here, $\gamma_{\nu}$, $S$, and $B_{qg}$ have the biggest post-fit impact on $\alpha_{s}(m_{Z})$ and they are also the ones that are somewhat constrained. This indicates that these parameters have the strongest correlation with $\alpha_{s}(m_{Z})$. Of course, to study the exact correlations between all parameters in detail, one has to examine the post-fit covariance matrix.
@@ 4.2.3 Profiling against N4LL
Figure 5: Left panel: Uncertainties in the $q_{T}$ spectrum at N2+1LL relative to the N4LL, before (yellow band) and after (orange band) profiling the TNPs. The different lines show the post-fit relative impact of each TNP. Right panel: Post-fit constraints on the TNPs (error bars) and their impact on $\alpha_{s}(m_{Z})$, with the solid (dashed) grey band showing the impact of the post-fit downward (upward) TNP variations. The stars indicate the true values of the TNPs.
@@
In figure 5, we consider the case where the theory model is given by the N2+1LL prediction. The left panel shows the pre-fit and post-fit uncertainties for the $q_{T}$ spectrum analogous to the left panel of figure 4 but now relative to the true N4LL result. The profiling again leads to a significant reduction in the uncertainty. Furthermore, we see that the post-fit prediction is now pulled toward the true result, with a significant adjustment of the shape of the spectrum. Whilst the shape of the pre-fit prediction differs from the true shape, the post-fit prediction closely reproduces it.
@@
The right panel of figure 5 shows the corresponding pull plot for the TNPs (analogous to the right panel of figure 4) with the true values of the TNPs marked by stars. The fitted TNPs at N2+1LL now exhibit strong pulls toward their true values, showing that, as for the $q_{T}$ spectrum, the fit is moving the theory model in the right direction also at the level of individual TNPs. This is exactly what we would like to happen in this situation where the data has a higher precision than the pre-fit theory constraint. However, some of the TNPs, such as $\gamma_{\nu}$, $B_{qq}$, and in particular $B_{qg}$, are quite strongly constrained now. A tight post-fit constraint implies that the corresponding parameter at the next higher-order becomes relevant. In other words, we are starting to “overfit” the N2+1LL theory model. The precision of the data is starting to exceed the inherent precision of the theory model, suggesting that it is becoming insufficient and using a higher-order theory-model might be warranted.
@@
Figure 6: Left panel: Uncertainties in the $q_{T}$ spectrum at N3+1LL relative to the N4LL, before (yellow band) and after (orange band) profiling the TNPs. The different lines show the post-fit relative impact of each TNP. Right panel: Post-fit constraints on the TNPs (error bars) and their impact on $\alpha_{s}(m_{Z})$, with the solid (dashed) grey band showing the impact of the post-fit downward (upward) TNP variations. The stars indicate the true values of the TNPs.
@@
In figure 6 we repeat the same exercise with the N3+1LL theory model. The post-fit uncertainties for the $q_{T}$ spectrum are very similar to those in figure 4 when profiling N3+1LL against itself, except that now the central value is shifted toward the true result. The same holds at the level of individual TNPs. Their post-fit uncertainties are very similar to before, while their central values are shifted toward their true values, which is again as we would like it to be. Compared to the N2+1LL case, the TNPs are now less constrained, as expected, with no indications yet of overfitting the theory model.
@@
Figure 7: Perturbative uncertainties on $\alpha_{s}(m_{Z})$ when profiling the TNPs at different orders against the central N4LL prediction. The numerically displayed uncertainties are in units of $10^{-3}$.
@@
In figure 7, we present the results for $\alpha_{s}(m_{Z})$ when profiling the TNPs at various orders against the N4LL pseudodata. The uncertainty on $\alpha_{s}(m_{Z})$ is only slightly larger at N2+1LL than at N3+1LL, which highlights the fact that after profiling it is largely driven by the constraints imposed by the measurement on the theory model. In addition, we also include results at N3+0LL and N4+0LL – these orders serve as approximations of N3+1LL and N4+1LL, respectively, see appendix B.2 for their definition and a more detailed discussion. While these approximations lack the correct formal structure of their corresponding higher orders, we include them here for completeness. Note that although N4+0LL is technically our currently highest available order, we prefer to use N3+1LL as our nominal theory model, because it has the correct theory correlation structure. Furthermore, its theory precision is sufficiently high compared to the measurement precision and after profiling it yields uncertainties comparable to N4+0LL.
@@
At both N2+1LL and N3+1LL we observe a small bias in the fitted central value of $\alpha_{s}(m_{Z})$. First note that when fitting the lower-order theory model against the higher-order pseudodata, we intentionally introduce a mismatch. Without profiling, this naturally leads to a potential bias in $\alpha_{s}(m_{Z})$ as it potentially absorbs some of the deficiencies of the theory model. This type of bias is actually what is estimated and supposed to be covered by the scanning approach (for either TNPs or scale variations). Indeed, fitting the central N2+1LL and N3+1LL models against N4LL and evaluating the uncertainty by scanning over TNPs, we find $\alpha_{s}(m_{Z})=0.1211\pm 0.0044$ and $\alpha_{s}(m_{Z})=0.1195\pm 0.0018$, respectively. The biases compared to the correct value of $0.118$ are thus indeed covered by their respective scanning uncertainties.